Find the indicated complement. Find P(A), given that P(A) = 0.932
Answer #1

The event A-   indicates the complement of event A.

Also it is given that p(A) = 0.932

so by using property of complementary events,

p(A- ) = 1- p(A) = 1-0.932 = 0.068

# Answer is 0.068